OpenStudy (anonymous):

Six letters are picked. Find the chance that they can be arranged to form the word RANDOM.

OpenStudy (anonymous):

Maybe this will get the answer you want...(I'm not very good at this stuff...) 1/26)*(1/25)*(1/24)=

OpenStudy (anonymous):

Yes, thank you

OpenStudy (anonymous):

total no. of letters in English alphabet=26 chances of forming a word RANDOM= 1/26*1/25*1/24*1/23*1/22*1/21=1/(26*25*24*23*22*21)